On a cold winter day, the inside of a window of a heated building is colder than the inside surrounding air. Natural convectiSolve the following problem using theories from heat transfer.
Answer #1

Given that

Answer - Given that A window of heating building in colder than the inside & wwounding air. Natural convection. (a) window as(0) window 1.5 m tall. 1 m wide Toroc Room temperature Tour emissivity 20695 Top 20°c. Noi = C. Ray Br. I Toret . = 273+Tos 36 Rayligh Number Rą = 9B (Teo-Ts) 23 --9.81*3.2713 X1020 via 7,426x105 x 1.944x100 . 8.15X1059 7181. So turbulent how cao-lo,- 3.27 X1.5X (26) +9.05x15 2439_2734)./ 8.67X16€. = 98.1 t. 13.5+ (2437_2.734) (5.67 ) x 10-45 24486 w Heat loss = 244.860 /